OpenStudy (anonymous):

Kate has 21 coins (nickels and dimes) in her purse. How many nickels and dimes does she have if she has $1.50?

OpenStudy (anonymous):

3 nickels and 18 dimes 5 nickels and 16 dimes 12 nickels and 9 dimes 15 nickels and 6 dimes

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

what do you think it is?

OpenStudy (anonymous):

idk. Thats why im asking the question.

OpenStudy (anonymous):

ok well dimes are worth 10 and nickels are worth 5 so in the answers look 3 nickels and 18 dimes just in dimes there is to much 18*10=180 5 nickels and 16 dimes same here 16*10=160 what 12 nickels and 9 dimes whats 12*5=____ 9*10=____ 15 nickels and 6 dimes whats 15*5=____ 6*10=____ thne just solve